A continuous wave train has 6 wave crests in a distance of 83.0 cm. If the source vibration
complete 14.0 cycles in 10.0 seconds, how far can one of these waves travel in 7.00 minutes?
1
Expert's answer
2020-06-23T12:58:46-0400
"(6-1)\lambda=0.83\to \lambda=0.166\ m"

"s=vt=\lambda f t=(0.166)(1.4)(7\cdot 60)=97.6\ m"